# Shahriar Noroozizadeh

### v2

@@ 3.1 Supervisory pressure does not explain geometric memory
(Explicit capacity pressures) Perhaps, the model memorizes geometrically because a lookup of local associations is impossible to represent. This restriction is either due to the parameter count or the design of the architecture, especially the embedding bottleneck.
@@
This logic can be dismantled as follows. First, we note that our architectures can express the associative model. Positive theoretical results in Nichani et al. [124] roughly prove that, with $m$ (frozen) embedding dimensions, and at least $m^{2}$ free parameters, an MLP layer can store $m^{2}$ many associations. Our models have $m\approx 400$ which means our large graphs can be stored associatively in our models.
@@
But maybe, there are yet other explicit pressures—such as normalizations and weight decay—that discourage associative storage. As a definitive refutation of such pressures, we make an empirical demonstration: there are experimental settings (with sufficient capacity and no bottleneck pressure) where geometric memory naturally arises; and in those very settings, an associative memory can be learned with the embeddings frozen and all else unchanged. Thus, there is demonstrably no explicit capacity pressure restricting an associative memory—and yet, a geometric one naturally emerges.
@@
(Associative memory can be artificially learned with our architectures) On various sequence models (Transformers, Mamba, neural networks) and various tiny graphs in Fig.˜1 and §C.3, a geometric memory arises naturally even though the same setup fits the data with frozen embeddings. [^3] This fit is purely associative since only one trainable layer exists in these architectures (see §B.2.2).
@@
Thus, explicit capacity pressures are not required to force a geometric memory in place. Maybe then the optimizer has a role to play. For instance, it is known that in classification settings, gradient descent takes longer to discover a lookup of artificially-randomized labels than natural labels [12]; it is also well-known that optimization tends to fit the data in progressively more complex ways [140, 184, 82, 12, 146, 19] and so a lookup would take relatively long to discover than other solutions. A similar effect can be proposed here:
@@
(Optimizational pressure) Perhaps, the model memorizes geometrically because an associative storage is harder (e.g., takes longer) to find by gradient descent.
@@
Refuting this, we argue that in sequence modeling, an associative lookup is significantly easier to find than the geometric one. In fact, during training, models typically first store the data associatively before storing geometrically. First, this is already (indirectly) evidenced at various points in literature (although not interpreted this way). In the graph tasks of [189, 174], the Transformer fits atomic facts well before it “groks” to learn circuits/representations that aid two-hop reasoning; in natural language [96] models memorize n-gram statistics well before representations become more organized, giving way to generalization.
@@
A more precise and vivid claim can be made here: Nichani et al. [124] observe that the associative lookup table $\sum_{(u,v)\in E}{\bf{\Phi}}(u)\otimes{\bf{\Phi}}(v)$ (from Section˜2.3) can be found by one step of gradient descent given a reasonably large embedding size on the correlation loss, $\sum_{(u,v)\in E}f(u)[v]$. [^4] This “one-step-lookup-table” phenomenon is distinctive of sequence modeling but not of a typical classification/regression task: for the lookup to be learned easily, we want the inputs near-orthogonal. This is true by design of the initial token embeddings of any sequence modeling task, but not necessarily true of the input distribution in classification/regression. We empirically verify this for the cross-entropy loss on our tiny graphs: associative memory forms in $2$ steps (Fig.˜7); whereas a geometric memory requires $100$ steps (Fig.˜8).
@@
Our intuition for this “associative first, geometry later” dynamics is that an associative memory merely stores the adjacency matrix $\bf{A}$} that is surficially available in the data; a geometric memory requires unearthing implicit information from the data, which requires more iterations—as we will see in §4, this computation corresponds to low-rank factorizing the adjacency matrix, or equivalently, computing a multi-hop adjacency matrix $\bf{A}$$^{\ell}$.
@@
(Associative memory is easier to find with gradient descent) Given sufficient embedding dimensionality and capacity, the associative storage can be discovered by full-batch gradient descent in as little as two steps on all our tiny graphs (Fig.˜7) while a geometric memory takes much longer to form (Fig.˜8). Similarly, we interpret the earlier findings of [189, 174, 96] as the data being fit associatively early during training, while a geometry (and therefore, reasoning) arising only much later.
Yet another plausible explanation for the geometry can be derived from the core of modern intuition about generalization, thanks to [122, 195]: gradient descent implicitly finds succinct fits of the data even without being explicitly told to do so. So, even if gradient descent is quick to discover associative memory, over time, perhaps it discovers a geometric organization as it gradually “compresses” the data. Indeed, such compressive or complexity-reducing effects have been brought up in recent analyses of geometry in language [96] and reasoning [174]. We translate this into our terminology:
@@
To refute this, we refute the underlying premise itself: counterintuitively, a geometric storage is not necessarily more succinct than an associative storage. To see why, we first clarify that unlike a typical classification task, or even a typical natural language task, our setting must be recognized as a memorization task (in the sense of Feldman [42]): the existence of an edge cannot be statistically surmised from the rest of the training set. Thus, we will show, for certain graphs, straightforward notions of succinctness do not break the tie between the two ways of representing the data.
Concretely, in generalization theories, the complexity of a lookup table is larger than the succinct one by a factor that scales polynomially with the number of datapoints. In contrast, in the path-star or cycle memorization task, this gap (measured through bit or norm complexity) can vanish:
(Geometric and associative memory are roughly equally succinct) For certain graphs, the complexity of the local associative memory both in terms of bits and $\ell_{2}$ norms is either equal to or at most twice those of the global geometric memory. [^5] Equality is achieved if there is no weight-tying or if only forward edges are stored. [^6] Note that a geometry appears even when storing only one direction of the edges in the smaller graphs where both forms of storage must be equally succinct; see §D.1.2.Informal proof and more explanations in §E.2.
Thus, while the familiar geometries of word embedding tasks [40, 130, 112] arise where a lookup is clearly more verbose, similar geometries arise even in tasks where redundancies are absent and a lookup may be just as succinct.
@@ 3.3 The memorization puzzle
The above arguments isolate a fundamental aspect of well-known neural geometries that cannot be easily explained by well-known learning pressures. Geometric representations emerge even in an atomic fact-memorization task that lacks statistical redundancies. Attempts to understand this through familiar intuitions from classification/regression tasks—from the lens of ease-of-optimization and succinctness—seem to fail us in the land of sequence models. [^7] Intuitively, this is because the geometric representations are of a fundamentally different nature from the well-studied feature representations in classification/regression tasks (see Remark 3). This gives rise to what we call the memorization puzzle:
@@
Answering this requires a more thoughtful argument, perhaps by borrowing other notions of complexity (e.g., flatness [84] or spectral norms [123] or something else not considered so far). One may also consider some other well-understood effects of depth and factorization in neural networks [74, 11]; although, we cast doubt even on these effects; see Remark˜1.
@@
Figure 7: Associative memory can be discovered quickly by gradient descent, given a sufficiently wide model, and a sufficiently large learning rate: For the various tiny graphs described in §C.3, and for our TinyNN model (with frozen embedding/unembedding layers and one wide trainable weight matrix to prevent a geometry from taking over; see §B.2.2), we report memorization over timesteps of training with full-batch gradient descent and constant learning rate of $0.1$. For each vertex, we compute what fraction of its $k$ neighbors appear in the top $k$ next token probabilities; this is averaged across vertices (where $k$ varies across vertices). This value, which quantifies associative memorization, reaches its maximum within 22 steps of gradient descent; compare this to the longer time for geometric memory to form in Fig.˜8. This is evidence for our Section˜3.2: ease-of-discovery does not dictate what memory is formed.
(a) Tiny Path-Star
(b) Tiny Grid
(c) Tiny Cycle
(d) Tiny Irregular
Figure 8: Geometric memorization takes much longer for gradient descent to discover: For the same architecture as in Fig.˜7 but with learnable embeddings, we roughly quantify geometric memorization over time via cosine distance across all pairs of vertices, alongside 2D visualizations. Observe that the geometries emerge after about a $200$ steps or more, compared to the 22 steps required for associative memory in Fig.˜7. This is evidence for Section˜3.2. While this is for learning rate $0.001$ (where the geometries are cleanest), the observation holds for larger learning rates too; please see Fig.˜22.
Setting aside the competition with the associative memory, we can still ask questions about the geometry: what is the geometry, how does it correspond to global information, and how does it arise without explicit pressures? Recall that we find these geometries not only in models with attention layers, but also in Mamba and in classical deep networks; to focus on the core phenomenon, we study the simplest yet non-trivial multi-layered gradient-descent-trained model: a 2-layer linear network consisting only of an embedding and unembedding layer (weight-tied), trained on the same loss (cross entropy) and the same dataset (of local edges). Equivalently, this is a one-layer, 1-hop Node2Vec model. Here, associative memory is architecturally prohibited, [^8] Node2Vec with weight-tying cannot implement associative memory at least in the form we care about. A more contrived form is possible; see §E.3. so we can pursue a clean analysis of the geometry alone.
4.1 Where does the geometry come from?
A precise characterization of what embeddings are learned even in such simple models is an open question. But a rich line of work—with key assumptions about various pressures outlined shortly—points to a low-rank spectral bias: the learned embeddings often align with the top (non-degenerate) eigenvectors of the negative graph Laplacian (roughly given by ${\bf{A}}-{\bf{D}}$, the difference between the adjacency and the diagonal degree matrix). Our figures corroborate this: the Node2Vec embeddings in Fig.˜1 right column match the top eigenvectors—called the Fiedler vector(s)—in Fig.˜9 left column.
This leads us to a few key insights. First, given the similarity in the geometries of Node2Vec and our deep sequence models (Transformer, Mamba and neural networks), such a spectral bias must indeed be at the heart of all such deep models, regardless of architectural quirks like self-attention and residual connections. However, the Node2Vec geometries turn out to be more well-organized than that which the Transformer exhibited (in Fig.˜1 middle column). Thus, we conjecture a similar spectral geometry in Transformers but—in hindsight—mildly adulterated by the competition with local associative memory.
Conjecture 1.
@@ 4.2 Where does the global information come from?
Another key insight is that the top eigenvectors, which are known to encode global structures in the graph, are the very source of global information that aided our reasoning task. Intuitively, the top eigenvectors of the adjacency matrix $\bf{A}$} are also the directions that become dominant in the $\ell$-hop adjacency matrix $\bf{A}$$^{\ell}$, directions that translate to global information. An alternative perspective is to view geometric memory as having factorized the adjacency matrix $\bf{A}$} while (softly) filtering out the lower eigenvectors; this extra computation has produced new information not surficially available in the adjacency data.
@@ 4.3 Where does the low-rank spectral bias come from?
While the geometry stems from a low-rank spectral bias, it is not clear where the bias itself comes from. Going back to Node2Vec theories (see §5.5), the literature suggests that Node2Vec-like models rely on the top eigenvectors under the assumptions of an explicit pressure of a bottleneck [93, 66, 165, 76, 152, 153, 14] or explicit regularization [83, 199] or an explicitly multi-hop supervision [137]. Our setting defies all these assumptions.
We hypothesize that these analyses inadvertently introduce explicit pressures in an attempt to simplify the dynamics e.g., nearly all analyses are on simpler-to-analyze losses such as the squared-error or a negative-sampling loss. On the other hand, the (more complex-to-analyze) cross-entropy loss is known to provide implicit pressures in certain other settings, such as in classification where it induces margin-maximization [162]; it is also the loss of choice in modern language models. Thus, we undertake an analysis of this loss, without assuming any explicit pressures. We emphasize that closed-form expressions here have remained an open question.
A priori, one would expect the unregularized cross-entropy-based Node2Vec system to behave in one of few ways: the system may diverge (to drive the probabilities to 0 or $1$), or may converge but only in direction (like in logistic regression [162]); the converged direction in turn, may be degenerate (e.g., node representations collapse to a single value) or not. Through empirical analysis, we suggest that none of these are true. Instead, the system appears to converge to a zero-gradient solution. It does so by elegantly achieving two properties, one of which is a low-rank spectral bias:
(How spectral bias emerges without typical pressures) In a $1$-layer, 11-hop Node2Vec model with the embedding ${\bf{V}}\in\mathbb{R}^{n\times m}$ of $n$ nodes, with the dynamics denoted as $\dot{{\bf{V}}}(t)=\eta{\bf{C}}(t){\bf{V}}(t)$ (where ${\bf{C}}(t)\in\mathbb{R}^{n\times n}$ is a time-dependent co-efficient matrix), the converged solution for our small graphs is such that (a) the columns of embedding matrix $\bf{V}$} span the graph’s Fiedler-like vectors [^9] We use “Fiedler-like” to denote not just the second-top eigenvectors of the negative Laplacian matrix, but also some of the subsequent top eigenvectors., and (b) the co-efficient matrix $\bf{C}$} has those same vectors in its null space (Fig.˜9). Crucially, this dynamic does not need a low rank constraint; the embedding size $m$ can be larger than the graph.
@@
Figure 9: Spectral geometry arises in Node2Vec without low-rank pressure (Section˜4.3) for tiny Path-Star, Grid, Cycle, and Irregular Graphs (top to bottom). (a) The Fiedler-like vectors of the graphs encode global structure; this structure mirrors the Node2Vec embeddings shown in Fig.˜1. (b; left) The evolution of eigenvector projections during training. (middle) The embedding matrix $\bf{V}$} contracts into space spanned by the Fiedler-like eigenvectors, evidenced by the projection norm ${\bf{V}}^{T}{\bf{e}}_{i}||_{2}$ converging to a stable, non-zero value; projections of other eigenvectors diminish towards zero. (b; right) Concurrently, the null space of the co-efficient matrix $\bf{C}$} subsumes the Fiedler-like eigenvectors, in that the norm ${\bf{C}}{\bf{e}}_{i}||_{2}$ converges to 0. This spectral bias arises without a low dimensional constraint assumed in literature (embedding size $m=100$, much larger than nodes in graph).
Why do the dynamics play out this way? We provide in §F an empirically-informed intuition (and a partial proof) for a “self-stabilizing” dynamic. The system gradually filters out lower eigenvectors from the embedding matrix $\bf{V}$}, while increasing their presence in the co-efficient matrix $\bf{C}$}, zeroing out the gradients ${\bf{C}}{\bf{V}}$ over time. We leave open a formal analysis of convergence, its rate, and whether this can be characterized into a form of max-margin-style bias.
We emphasize that this two-layer analysis does not divulge anything about the competition between associative and geometric memories. That requires a study of deeper models, where the dynamics of different spectral components begin cross-interfering. What this two-layer analysis does is make progress on an open question for the simplest model possible. It also gives us an idea of how and what global information can arise naturally out of local supervision, devoid of any supervisory, bottleneck-driven, or regularizing pressure. These spectral dynamics must be central to any gradient-descent-trained deep sequence model.

### v3

@@ 3.2 Capacity or optimizer pressures do not explain geometric memory
(Explicit capacity pressures) Perhaps, the model memorizes geometrically because a lookup of local associations is impossible to represent, either due to the parameter count or the design of the architecture, especially the embedding bottleneck.
@@
This logic can be dismantled as follows. First, our architectures can express the associative lookup. Positive theoretical results in Nichani et al. [133] roughly prove that, with $m$ (frozen) embedding dimensions, and at least $m^{2}$ free parameters, an MLP layer can store $m^{2}$ many associations. Our models have $m\approx 400$ which means our large graphs can be stored associatively in our models.
@@
But maybe, there are yet other explicit pressures such as dropout and weight decay that discourage associative storage. As a definitive refutation of such pressures, we make two empirical demonstrations. First, we show that with those pressures switched off, a geometry still arises in high-model-width settings. Second, in the very settings where geometries emerge, an associative memory can be learned with the embeddings frozen and all else unchanged. Thus, there is demonstrably no explicit capacity pressure restricting an associative memory—and yet, a geometric one naturally emerges.
@@
(Associative memory can be artificially learned with our architectures) First, a geometry arises even without the pressures of weight decay or dropout (§C.3.1). Next, in various sequence models (Transformers, Mamba, neural networks) and various tiny graphs in Fig.˜1 and §C.3, a geometric memory arises naturally even though the same training setup can fit the data associatively, with embeddings frozen and one trainable intermediate layer. [^2] This fit is purely associative since only one trainable layer exists in these architectures (see §B.2.2).
@@
Thus, explicit capacity pressures are not required to force a geometric memory in place. Maybe then the optimizer has a role to play. For instance, it is known that in classification, gradient descent takes longer to discover a brute-force lookup [13]; it is also well-known that optimization tends to fit the data in progressively more complex ways [151; 195; 88; 13; 19; 21] and so a lookup would take relatively long to discover than other solutions. A similar effect can be proposed here:
@@
(Optimizational pressure) Perhaps, the model memorizes geometrically because an associative storage takes longer to find by gradient descent.
(a) Associative Memorization
(b) Geometric Memorization
Figure 6: Associative first, geometric later: Associative memory (left) is only two steps away under gradient descent whereas geometric memory (right) takes much longer to form for the tiny path-star graph. More details and similar plots for various other graphs and learning rates are in Figs.˜21 and $2$.
@@
Refuting this, we argue that in sequence modeling, an associative lookup is significantly easier to find than the geometric one. In fact, during training, models typically first store the data associatively before storing it geometrically. First, this is already (indirectly) evidenced at various points in literature (although not interpreted this way). In the graph tasks of [200; 184], the Transformer fits atomic facts well before it “groks” to learn circuits/representations that aid two-hop reasoning; in natural language [103] models memorize n-gram statistics well before representations become more organized, giving way to generalization. Indeed, we propose our setting as arguably the most minimal instance of grokking [145].
@@
A more precise and vivid claim can be made here: Nichani et al. [133] observe that the associative lookup table $\sum_{(u,v)\in E}{\bf{\Phi}}(u)\otimes{\bf{\Phi}}(v)$ (from Section˜2.3) can be found by one step of gradient descent given a reasonably large embedding size on the correlation loss, $\sum_{(u,v)\in E}f(u)[v]$. [^3] This “one-step-lookup-table” phenomenon is distinctive of sequence modeling but not of a typical classification/regression task: for the lookup to be learned easily, we want the inputs near-orthogonal. This is true by design of the initial token embeddings of any sequence modeling task, but not necessarily true of the input distribution in classification/regression.
@@
We empirically verify this in Fig.˜6 for the cross-entropy loss on our tiny graphs: for wide models, associative memory forms in $2$ steps, whereas a geometric memory requires $100$ steps. Our intuition is that an associative memory is easy to find for it merely stores the adjacency matrix $\bf{A}$} surficially available in the data; a geometric memory requires unearthing implicit information from the data, which requires more iterations—as we will see in §4, this computation corresponds to low-rank factorizing the adjacency matrix, or equivalently, computing a multi-hop adjacency matrix $\bf{A}$$^{\ell}$.
@@
(Associative memory is easier to find with gradient descent) Given sufficient embedding dimensionality and capacity, associative storage can be discovered in as little as two steps on all our tiny graphs (Figs.˜6(a) and 21) while a geometric memory takes much longer to form (Figs.˜6(b) and 22). Similarly, we interpret the earlier findings of [200; 184; 103] as the data being fit associatively early during training, while a geometry (and therefore, reasoning) arising only much later.
Yet another plausible explanation for the geometry can be derived from modern intuition about generalization [131; 207]: gradient descent implicitly finds succinct fits of the data even without being explicitly told to do so, and thus evades a complex lookup. So, even if gradient descent is quick to discover associative memory, over time, perhaps it discovers a geometric organization as it gradually “compresses” the data. Indeed, such compressive or complexity-reducing effects have been brought up in recent analyses of geometry in language [103] and reasoning [184]. We translate this into our terminology:
@@
To refute this, we challenge the underlying premise itself. Counterintuitively, a geometric storage is not necessarily more succinct than an associative storage. To see why, we first clarify that unlike a typical classification task, or even a typical natural language task, our setting must be recognized as a memorization task (in the sense of Feldman [46]): the existence of an edge cannot be statistically surmised from the rest of the training set. Thus, straightforward notions of succinctness do not break the tie between the two ways of representing the data, at least for sparse graphs. Concretely, in the path-star or cycle memorization task, we find no complexity gap, when measured through bit or norm complexity—contrast this against the generalization setting, where the complexity of a lookup table would be significantly larger than the succinct one, typically by a factor that scales polynomially with the number of datapoints.
(Geometric and associative memory are equally succinct) For sparse graphs where the vertex and edge counts scale similarly (e.g., path-star or cycle), the complexity of associative memory both in terms of bits and $\ell_{2}$ norms is either equal to or at most twice those of geometric memory. [^4] Equality is achieved if there is no weight-tying or if only forward edges are stored. Informal proof and more explanations in §E.2.
Thus, while geometries are well-known in word embedding tasks [44; 140; 121] where a lookup is clearly more verbose, this result isolates what is not obvious about them: geometries arise even when redundancies are absent and a lookup may be just as succinct.
@@ 3.3 The memorization puzzle
The above arguments isolate a fundamental aspect of well-known neural geometries that cannot be easily explained by well-known learning pressures. Geometric representations emerge even in an atomic fact-memorization task that lacks statistical redundancies, and attempts to understand this through familiar intuitions from classification/regression tasks—from the lens of ease-of-optimization and succinctness—seem to fail us in the land of sequence models. [^5] Intuitively, this is because the geometric representations are of a fundamentally different nature from the well-studied feature representations in classification/regression tasks (see Remark 3). This gives rise to what we call the memorization puzzle:
@@
Answering this requires a more thoughtful argument, perhaps by borrowing other notions of complexity (e.g., flatness [90] or spectral norms [132] or something else not considered so far), or by invoking other well-understood effects of depth and factorization in neural networks [80; 12]—although, we cast doubt even on these effects; see Remark˜1—or by invoking certain rank-minimizing and feature-sparsifying effects [190; 52; 205; 30; 11].
Setting aside the competition with the associative memory, we can still ask questions about the geometry: what is the geometry, how does it correspond to global information, and how does it arise without explicit pressures? Recall that we find geometries not only in models with attention layers, but also in Mamba and in classical deep networks; to focus on the core phenomenon, we study the simplest yet non-trivial multi-layered gradient-descent-trained model: a 2-layer linear network consisting only of an embedding and unembedding layer (weight-tied), trained on the same loss (cross entropy) and the same dataset (of local edges). Equivalently, this is a 1-hop Node2Vec model with a 1-layer embedder. Here, associative memory is architecturally prohibited, [^6] Node2Vec with weight-tying cannot implement associative memory at least in the form we care about. A more contrived form is possible; see §E.3. so we can pursue a clean analysis of the geometry alone.
4.1 Where does the geometry come from?
A precise characterization of what embeddings are learned even in such simple models is an open question. But a rich line of work—with key assumptions about various pressures outlined shortly—points to a low-rank spectral bias of gradient descent: the learned embeddings align with the top (non-degenerate) eigenvectors of the negative graph Laplacian (roughly given by ${\bf{A}}-{\bf{D}}$, the difference between the adjacency and the diagonal degree matrix). Our figures corroborate this: the Node2Vec embeddings in Fig.˜1 right column match the top eigenvectors—called the Fiedler vector(s)—in Fig.˜8 left column.
There are two notable insights from comparing the geometries of Node2Vec against our various deeper sequence models. First, these two geometries are similar, suggesting that the known spectral bias in 2-layer models must indeed be at the heart of all such deep models, regardless of architectural quirks. But notably, the Node2Vec geometries are much more well-organized than what the Transformer exhibits (in Fig.˜1 middle column). In total, we conjecture a similar spectral geometry in Transformers but—in hindsight—mildly adulterated by the competition with local associative memory.
Conjecture 1.
@@ 4.1 Where does the geometry come from?
Figure 7: Low-rank spectral bias arises naturally in 22-layer cross-entropy-trained models even without explicit pressures. We consider a wide model ($d_{\text{embedding}}=100$, much larger than nodes in graph, so no rank constraint) trained on local supervision. We report the projection of the weights along the eigenvectors (normalized by that of the top eigenvector). As stated in Section˜4.3, the model is naturally skewed towards the top vectors, in fact even more skewed than the eigenvalues themselves e.g., the model learns only a 2-dimensional embedding of cycle and grid graphs. Note that this observation is specific to the skew within the positive eigenvectors; the absence of negative directions is simply because the weight-tied Node2Vec can only express positive semi-definite matrices.
@@ 4.2 Where does the global information come from?
The spectral bias also tells us where the global information comes from: the top eigenvectors, which are known to encode global structures in the graph. Intuitively, the top eigenvectors of the adjacency matrix $\bf{A}$} are also the directions that become dominant in the $\ell$-hop adjacency matrix $\bf{A}$$^{\ell}$, directions that translate to global information. An alternative perspective is to view geometric memory as having factorized the adjacency matrix $\bf{A}$} while (softly) filtering out the lower eigenvectors; this extra computation has produced new information not surficially available in the adjacency data.
@@ 4.3 Where does the low-rank spectral bias come from?
While the geometry stems from a low-rank spectral bias, we argue it is not clear where the bias itself comes from. Going back to Node2Vec theories (see §5.5), the literature suggests that Node2Vec-like models rely on the top eigenvectors by assuming the explicit pressure of a bottleneck [100; 72; 175; 82; 162; 163; 15] or explicit regularization [89; 211] or an explicitly multi-hop supervision [148]. Our setting defies all these assumptions:
(Spectral bias without explicit pressures) A 2-layer Node2Vec-style model under cross-entropy loss naturally exhibits a low-rank spectral bias even without explicit pressures (i.e., with large width, no regularization, and 1-hop supervision) as in Fig.˜7.  [^7] We emphasize that 2-layer models show a particularly strong low-rank behavior not seen in deeper models. For instance, compare the two-layer cycle graph model in Fig. 7 with the three-layer ones in §C.3.1.
Prior analyses inadvertently introduce explicit pressures in an attempt to simplify the dynamics under other losses, e.g., nearly all analyses are on simpler-to-analyze losses such as the squared-error or a negative-sampling loss. On the other hand, the (more complex-to-analyze) cross-entropy loss is known to provide implicit pressures in certain other settings, such as in classification where it induces margin-maximization [172]; it is also the loss of choice in modern language models. Thus, our subsequent interest is in analyzing this loss, without assuming any explicit pressures, with the hope of explaining Section˜4.3.
A priori, one would expect the unregularized cross-entropy-based Node2Vec system to behave in one of a few ways: the system may diverge (to drive the probabilities to 0 or $1$), or may converge but only in direction (like in logistic regression [172]); the converged direction in turn, may be degenerate (e.g., node representations collapse to a single value) or not. Through empirical analysis, we suggest that none of these are true. Instead, the system appears to converge to a zero-gradient solution. It does so by elegantly achieving two properties, one of which is a low-rank spectral bias:
(How spectral bias emerges without typical pressures) In a 2-layer Node2Vec-style model with the embedding ${\bf{V}}\in\mathbb{R}^{n\times m}$ of $n$ nodes, with the dynamics denoted as $\dot{{\bf{V}}}(t)=\eta{\bf{C}}(t){\bf{V}}(t)$ (where ${\bf{C}}(t)\in\mathbb{R}^{n\times n}$ is a time-dependent co-efficient matrix), the converged solution for our small graphs is such that (a) the columns of embedding matrix $\bf{V}$} span the graph’s Fiedler-like vectors [^8] We use “Fiedler-like” to denote not just the second-top eigenvectors of the negative Laplacian matrix, but also some of the subsequent top eigenvectors., and (b) the co-efficient matrix $\bf{C}$} has those same vectors in its null space (Fig.˜8). Crucially, this dynamic does not need a low rank constraint; the embedding size $m$ can be larger than the graph.
@@
Figure 8: How a spectral geometry arises in Node2Vec without low-rank pressure (Section˜4.3) for various graphs (top to bottom). (a) The Fiedler-like vectors of the graphs encode global structure; this structure mirrors the Node2Vec embeddings shown in Fig.˜1. (b; left) The evolution of eigenvector projections during training. (middle) The embedding matrix $\bf{V}$} contracts into space spanned by the Fiedler-like eigenvectors, evidenced by the projection norm ${\bf{V}}^{T}{\bf{e}}_{i}||_{2}$ converging to a stable, non-zero value; projections of other eigenvectors diminish towards zero. (b; right) Concurrently, the null space of the co-efficient matrix $\bf{C}$} subsumes the Fiedler-like eigenvectors, in that the norm ${\bf{C}}{\bf{e}}_{i}||_{2}$ converges to 0 (or to a small value). This spectral bias arises without a bottleneck assumed in literature (embedding size $m=100$, much larger than nodes in graph).
Why do the dynamics play out this way? We provide in §F an empirically-informed intuition (and a partial proof) for a “self-stabilizing” dynamic. The system gradually filters out lower eigenvectors from the embedding matrix $\bf{V}$}, while increasing their presence in the co-efficient matrix $\bf{C}$}. Together, this zeroes out the gradients ${\bf{C}}{\bf{V}}$ over time. We leave open a formal analysis of what exactly the low-dimensional model converges to, the convergence rate, and whether this can be characterized into a form of max-margin-style bias:
Open Question 2: Low-rank Bias of 2-Layer Cross-Entropy Model Why (and under what conditions) does a wide, 2-layer, weight-tied model, trained on the cross-entropy loss over edge bigrams from a graph dataset, converge to only a few top eigenvectors of the adjacency matrix even in the absence of explicit pressures? What exactly is the solution learned and what is the rate of convergence?
We emphasize that a two-layer analysis (including our partial analysis) does not divulge anything about the competition between associative and geometric memories. That requires a study of deeper models, where the dynamics of different spectral components begin cross-interfering. What our own two-layer analysis does is make progress on an open question for the simplest model possible. It also gives us an idea of how and what global information can arise naturally out of local supervision, devoid of any supervisory, bottleneck-driven, or regularizing pressure. These spectral dynamics must be central to any gradient-descent-trained deep sequence model.